# Beyond the Dashboard: Investigating Distracted Driver Communication Preferences for ADAS

### v1

@@ Acknowledgements
Figure 6: Participant responses for the frequency with which they perform non-driving related activities while driving.
-C Demographics Questionnaire
@@
-D Driving Experience Questionnaire
@@
We also asked participants about the frequency with which they drive with the following distractions: Alcohol, Recreations Drugs, Sleepiness/Drowsiness, Text messages, Phone calls, Video calls, Secondary activities such as applying make up, eating, etc, Passengers, Highway hypnosis.
The aggregated responses for the frequency with which participants perform non-driving related activities are shown in Figure 6. Overall, participants attempt to keep their sources of distraction at a minimum. Taking phone calls while driving is reported as the most plausible source of distraction, with drowsiness and other passengers in the vehicle following suit. This order reinforces our prior belief about the sources of distraction and how to capture them within cars using audio-visual recordings as discussed in Section IV-C.

### v2

@@ Acknowledgements.
Appendix A Appendix
Figure 6. Participant responses for the frequency with which they perform non-driving related activities while driving.
[…]
[The responses of participants about which factors might distract them.]A Boxplot showing that phone calls, passengers, and drowsiness are the largest contributors towards driver distraction
A.1. Demographics Questionnaire
@@
A.2. Driving Experience Questionnaire
@@
We also asked participants about the frequency with which they drive with the following distractions: Alcohol, Recreations Drugs, Sleepiness/Drowsiness, Text messages, Phone calls, Video calls, Secondary activities such as applying make-up, eating, etc, Passengers, Highway hypnosis.
The aggregated responses for the frequency with which participants perform non-driving related activities are shown in Figure 6. Overall, participants attempt to keep their sources of distraction at a minimum. Taking phone calls while driving is reported as the most plausible source of distraction, with drowsiness and other passengers in the vehicle following suit. This order reinforces our prior belief about the sources of distraction and how to capture them within cars using audio-visual recordings as discussed in Section 3.3.